OpenStudy (anonymous):

Find a ordered pair for... 5x+y = 9

OpenStudy (phi):

if you make x = 0, in your equation, what do you get for y ?

OpenStudy (anonymous):

1,4

OpenStudy (anonymous):

1,4

OpenStudy (anonymous):

1,4

OpenStudy (anonymous):

5x+y=9 I will try to solve for y. 5x+y=9 5x+y+−5x=9+−5x(Add -5x to both sides) y=−5x+9 y=−5x+9

OpenStudy (anonymous):

So yeah, 1,4